OpenStudy (anonymous):

How many solutions are there for the system shown below? x^2+4y^2=100 4y-x^2=-20 A. 4 B. 3 C. 1 D. 2

OpenStudy (callisto):

x^2+4y^2=100 -(1) 4y-x^2=-20 => x^2 = 4y+20 - (2) Put (2) into (1) and solve for y, see how many solution(s) you can get Then solve x using the value of y, and determine the number of solution(s) of x Can you do it now?

OpenStudy (anonymous):

4y^2 + x^2 + 4y - x^2 = 100 - 20 4y^2 + 4y = 80 y^2 + y = 20 y^2 + y - 20 = 0 y = (-1 +/- sqrt(1 + 80)) / 2 y = (-1 +/- 9) / 2 y = -10/2 , 8/2 y = -5 , 4 4y - x^2 = -20 4 * (-5) - x^2 = -20 -20 - x^2 = -20 0 = x^2 x = 0 4 * 4 - x^2 = -20 16 - x^2 = -20 36 = x^2 x = -6 , 6 Solutions: (0 , -5) (-6 , 4) (6 , 4)